OpenStudy (anonymous):

In which quadrant does the solution of the system fall? y = x – 1 y = –3x – 5

hero (hero):

y = y x - 1 = -(3x+5) 1-x = 3x+5 4x = -4 x = -1 y = 2 (-1,2) second quad